A golf ball is hit into the air, and its height y in feet after x seconds is given by y=-17x^2+238x
a)find the vertex of this parabola
b) intercept the vertex

Answer by stanbon(75887)

A golf ball is hit into the air, and its height y in feet after x seconds is given by y=-17x^2+238x
a)find the vertex of this parabola
vertex occurs when x = -b/(2a) = -238/(2*-17) = 7
Find f(7) = -17*49+238*7 = 833
Vertex:: (7,833)
----------------------
b) intercept the vertex
What do you mean by this?
Cheers,
Stan H.
---------------